OpenStudy (anonymous):

Suppose you are taking a true-false test with twelve questions. If you guess at the answers on this test, find the probability of getting exactly eight correct answers

OpenStudy (anonymous):

1/2 chance of getting 100

OpenStudy (anonymous):

Exactly eight correct answers

OpenStudy (anonymous):

forgot to finish question-lol

OpenStudy (anonymous):

im not sure about it I think it would be 8/12 cuz you got 2 correct which is also 2/3

OpenStudy (anonymous):

cuz you got 8*

OpenStudy (anonymous):

how would I round that to three decimal points?

OpenStudy (anonymous):

0.08

OpenStudy (raden):

number events of (T T T T T T T T F F F F) = 12!/8!4! = ... it is a start

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Direct application of binomial distribution with n=12, p=1/2. Plug in the numbers yourself!

OpenStudy (raden):

then we knowed that the probability if the correct answer also incorrect answer is 1/2 resfectively, and assumed this case is independet event. so, the probability of P(8T's and 4F's) = (1/2)^12 * 12!/8!4! = ...

OpenStudy (raden):

@MistrisKat , did u get it what's the value of 12!/8!4! = ...

OpenStudy (raden):

do u know about factorial ?

OpenStudy (anonymous):

not really :(

OpenStudy (anonymous):

am I multiplying 12 x 8x4?

OpenStudy (anonymous):

not sure what the ! is doing in there, what it's function?

OpenStudy (anonymous):

factorials are like 8.7.6.5.4.3.2.1. right?

OpenStudy (raden):

yep, that's right so, what's the value of 

OpenStudy (anonymous):

working on it just downloaded a fractional calculator..

OpenStudy (anonymous):

12=479001600

OpenStudy (anonymous):

not sure how to do bottom piece. separately 8=40320 and 4 =24

OpenStudy (raden):

easier if u doing cancel out for the similar term :